In a chemistry class, 12 L of a 12% alcohol solution must be mixed with 20% solution to get a 14% solution. how many liters of the 20% solution are needed
Answer by mananth(16949)

percent ---------------- quantity
Alcohol I 12 ---------------- 12 gallons
Alcohol II 20 ---------------- x gallons
Mixture 14 ---------------- 12 + x gallons

12*12+20x= 14 ( 12 + x )
144 + 20 x = 168 + 14 x
20 x + -14 x = 168 - -144
6 x = 24
/ 6
x = 4
4 gallons of 20 % Alcohol II